# An Online Learning Approach to Prompt-based Selection of Generative Models

### v2

@@ Appendix D Additional Experimental Details and Results
We provide more examples showing that prompt-based generative models can outperform for text prompts from certain categories while underperforming for other text categories (see Figures 3, 4, and 6).
@@
Figure 4: Prompt-based generated images from UniDiffuser (Bao et al.,, 2023) and PixArt-$\alpha$: UniDiffuser attains a higher CLIPScore in generating type “train” prompts (35.29 versus 34.25) while underperforms for type “baseball bat” prompts (32.51 versus 34.30).
@@
4. Implementation details.
We use the CLIP-based features of the prompts as the context vector (Cherti et al.,, 2023) for tasks of T2I and T2V generation, and we use the CLIP-based features of the images in the task of image captioning. We set both the exploration and regularization parameters $\alpha,\eta=1$ in all the experiments. Two hyperparameters have to be chosen. The first one is the parameter $\gamma$ in the polynomial and radial basis function (RBF) kernels, which are given by
$k_{\textup{poly3}}^{\gamma}(x_{1},x_{2})=(\gamma\cdot x_{1}^{\top}x_{2}+1)^{3},\quad k_{\textup{RBF}}^{\gamma}(x_{1},x_{2})=\exp(-\gamma\cdot\|x_{1}-x_{2}\|^{2}).$
In the experiments, we select $\gamma$ to be 5 and 1 for the polynomial and RBF kernel functions, respectively. The second hyperparameter is the number of random features in the RFF-UCB algorithm. In addition, the features are generated once for each sample and stored to save the computation of the RFF-UCB algorithm.
5. Ablation study on hyperparameters.
[…]
6. Comparison of running time.
We compare the running time of PAK-UCB and RFF-UCB, and the results are summarized in Figure 8.
Figure 8: Running time: The execution time of PAK-UCB-RBF (PAK-UCB using the RBF kernel) and RFF-UCB on Setup 4. PAK-UCB-RBF takes around 10 minutes to finish 2,000 iterations of model selection, while RFF-UCB uses less than 2 minutes. Results are averaged over 20 trials.
7. Results on Other Conditional Generation Tasks

### v3

@@ C.1 Results on Setups 1 and 2
We provide more examples showing that prompt-based generative models can outperform for text prompts from certain categories while underperforming for other text categories (see Figures 5, 7, and 9).
@@
Figure 7: Prompt-based generated images from UniDiffuser (Bao et al., 2023a, ) and PixArt-$\alpha$: UniDiffuser attains a higher CLIPScore in generating type “train” prompts (35.29 versus 34.25) while underperforms for type “baseball bat” prompts (32.51 versus 34.30).
@@
C.2 Results on LLM Selection
Figure 11: Sample prompt for random Sudoku-solving task (Setup 2)
Figure 13: Sample prompt for Python code completion task (Setup 2)
Figure 14: Sample prompt for C++-to-Java code translation task (Setup 2)
C.3 Results on Other Conditional Generation Tasks
@@
C.4 Implementation Details and Ablation Study
1. Hyperparameters and Implementation.
For the PAK-UCB-poly3 algorithm, we utilize the $k_{\textup{poly3}}^{\gamma}(x_{1},x_{2})=(1+\gamma\cdot x_{1}^{\top}x_{2})^{3}$ kernel with $\gamma=5.0,10.0$. For RFF-UCB-RBF, we utilize the $k_{\textup{RBF}}^{\sigma}(x_{1},x_{2})=\exp(-\frac{1}{2\sigma^{2}}\|x_{1}-x_{2}\|_{2}^{2})$ kernel with $200$-dimensional RFF (i.e., $D=200$), where we set $\gamma=0.5$ for LLM-related experiments (Setup 2) and $\gamma=5.0$ for the rest of experiments. In both algorithms, the regularization parameter is set to be $\alpha=1$. In addition, we set the exploration parameter $\eta=\sqrt{2\log(2G/\delta)}$, which follows our regret analysis in Sections 5.1 and 5.2. We choose the standard $\delta=0.05$.
2. Ablation study on hyperparameters.
We conduct ablation studies on the selections of parameter $\sigma$ in the RBF kernel function and the number of features in RFF-UCB-RBF. The results are summarized in Figures LABEL:fig:abl-gamma and LABEL:fig:abl-feat, respectively. We select $\sigma=1,3,5$, and $7$, and the number of features varying between 25, 50, 75, and 100. Results show that the RFF-UCB-RBF algorithm can attain consistent performance. Additionally, we test the PAK-UCB-poly3 algorithm with $\gamma=1,3,5$, and $7$ in the polynomial kernel and regularization parameter $\alpha=0.5,1.0$, and $1.5$. The results are summarized in Figures LABEL:fig:abl-poly-gamma and LABEL:fig:abl-poly-alpha, respectively.
3. Comparison of PAK-UCB-RBF and RFF-UCB-RBF.
We compare the performance of PAK-UCB-RBF and its RFF counterpart, i.e., the RFF-UCB-RBF algorithm. The results show that the RFF method can attains a similar performance with the original PAK-UCB-RBF algorithm (Figures LABEL:fig:abl-rff-1 and LABEL:fig:abl-rff-2). Moreover, the RFF-UCB-RBF algorithm can significantly speed up the computation (Figure 20).
Figure 20: Running time: The execution time of PAK-UCB-RBF (PAK-UCB using the RBF kernel) and RFF-UCB-RBF on Setup 4. PAK-UCB-RBF takes around 10 minutes to finish 2,000 iterations of model selection, while RFF-UCB-RBF uses less than 2 minutes. Results are averaged over 20 trials.